OpenStudy (anonymous):

solve x-7>10 x>3 x>7 x>17 x>70 solve 3x<21 there is supposed to be a line under the < on this one.

OpenStudy (anonymous):

for x-7>10 just solve it like a normal equation. So x-7 >10 (dont care about less greater than sign get x alone so, 10+7 = 17 therefore, x >17

OpenStudy (anonymous):

thanks! how about the other one?

OpenStudy (anonymous):

@JackleBee

OpenStudy (igreen):

3x < 21 Divide 3 to both sides of the equation: 21 / 3 = 7 x < 7

OpenStudy (anonymous):

ahh okay so do the same drill. 3x<21 So the less than or equal to sign doesn't do much. just do 21/3 to isolate x which is 7 so x = 7

OpenStudy (anonymous):

which is actually remember the signs x < 7

OpenStudy (anonymous):

if you were diving by negative the sign would flip * just saying